Determine the value of k for which the system of linear equations has no solution.
2x - y = -15
8x + ky = 14

Answer by shweta(56)

Given
2x-y=-15
8x+ky=14
They have no solution ,so
2/8=-1/k
1/4=-1/k
k=-4